# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ XV.2 Machine Learning Based Approaches
The ongoing revolution in machine learning has already had a significant impact in the study of many-body quantum physics Carleo and Troyer (2017); Gray et al. (2018); Vicentini (2021). These range from identification of various exotic phases of matter in hitherto analytically challenging models Carrasquilla and Melko (2017); Greplova et al. (2020), to Hamiltonian learning Wang et al. (2017), and investigation of foundational issues like Bell-nonlocality Deng (2018). Machine-learning algorithms, which learn from previously acquired data, also assist in solving optimization problems that arise in quantum sensing. For a general overview of machine-learning assisted quantum metrology, we refer the interested reader to the recent review by Huang et al Huang et al. (2024). As we describe below, quantum sensing has already been shown as an especially fruitful area of application for machine learning methods.
@@ XV.2.2 Quantum machine learning strategies
XVI Summary and Outlook
Sensing Strategy Pros Cons
Inteferometric Quantum Sensors
External parameter is
induced by phase shift
$U(\theta)=e^{i\theta G}$
1. Heisenberg scaling possible
2. Universal optimal measurement
1. Difficult preparation
2. Sensitive to decoherence and particle loss
3. Works only for phase shift operations
Many-Body Quantum Sensors at Equilibrium
Ground or Thermal
State based for Hamiltonians
of the form $H{=}H_{0}{+}H_{1}(\theta)$
1. Quantum enhanced sensitivity achievable
by gap-closing phase transitions
2. Super-Heisenberg scaling possible
3. Robust against imperfections
1. Fine tuning is needed (local sensing)
2. Optimal measurement parameter-value
dependent
3. Preparation might be resource consuming
Non-Equilibrium Many-Body Quantum Sensors
Non-Equilibrium Steady
State based
1. Initialization does not matter
2. Super-Heisenberg scaling possible
3. Works even with partial accessibility
4. Dissipation can also be harnessed
1. Fine tuning is needed (local sensing)
2. Optimal measurement parameter-value
dependent
3. Reaching steady state is time consuming
Transient State based
1. Easy initialization
2. Super-Heisenberg scaling possible
3. Time control for better precision
4. May operate across the entire phase
(no fine tuning)
5. Sequential measurements are adoptable
1. Decoherence is a limiting factor
2. Optimal measurement parameter-value
dependent
@@ 2. Optimal measurement parameter-value
Exploiting quantum features, such as entanglement, can enhance the precision of a sensor well-beyond the capacity of a classical sensor, the so-called quantum-enhanced sensing. Such enhanced in precision has placed quantum sensing as one of the pillars of emerging quantum technologies. In addition to such enhancement, the size versatility of quantum probes allows for achieving extremely high spatial resolution. The potential applications are immense covering from biological monitoring to mining, atomic clocks, navigation and space exploration.
@@
Interferometric quantum sensing has been the first demonstration of entanglement as a resource for sensing purposes. These types of sensors are very susceptible to decoherence and any interaction between particles is detrimental for their precision. An alternative approach is to exploit quantum many-body systems for sensing. Unlike interferometry-based quantum sensors, in many-body probes interaction plays a crucial role. In this review, we have explored different aspects of quantum sensing using many-body systems.
@@
Quantum many-body sensors have been used in two scenarios, namely equilibrium and non-equilibrium. In the former, quantum criticality has been identified as a resource for achieving quantum enhanced sensitivity. There are several types of quantum criticalities, each with their own characteristics. However, in order to achieve quantum enhanced sensitivity the criticality has to be accompanied by some sort of gap closing, for instance, Hamiltonian or Liouvillian spectral gap closing. This remarkable feature can be a useful guideline for identifying potential quantum sensors in new materials, which might be even designed by artificial intelligence. Although criticality-based quantum sensors offer several advantages, they also come with some drawbacks. First, the region over which quantum enhanced precision is achievable can be very narrow, requiring probe’s fine tuning. Second, preparing the probe near its critical point can be very resource-consuming because the preparation process slows down significantly as it approaches criticality. In order to avoid these drawbacks, non-equilibrium quantum probes have also been proposed. In such systems quantum superposition and entanglement can be generated during the dynamics and thus complex initialization is not needed. Moreover, in non-equilibrium quantum probes, the evolution time acts as an additional parameter that can be used as a resource to enhance precision. In fact, the precision is often enhanced monotonically as time goes on. Nonetheless, unlike criticality-based sensors, in non-equilibrium probes the criteria for achieving quantum enhanced sensitivity with respect to resources, e.g. system size, is not well-characterized. We summarize the advantages and disadvantages of each approach in Table 2.
@@
Despite current progress on many-body quantum sensing, there are still several open problems to explore. While criticality has been identified as a resource for sensing in equilibrium scenarios, the criteria which results in quantum enhanced sensing is not specified in non-equilibrium probes. Another problem which requires further investigation is the performance of quantum sensors under imperfect situations, such as the presence of decoherence. In general terms, while quantum sensors are naturally expected to be sensitive to the desired parameters, we would like them to be robust against other system’s parameters or noisy imperfections. Despite several works on noisy quantum metrology Escher et al. (2011b); Kołodyński and Demkowicz-Dobrzański (2013); Rossi et al. (2020); Demkowicz-Dobrzański and Maccone (2014); Demkowicz-Dobrzański et al. (2017); Dür et al. (2014); Nichols et al. (2016); Chaves et al. (2013); Alushi et al. (2024); Faist et al. (2023); Reiter et al. (2017); Mihailescu et al. (2023, 2024a), the notion of robustness has not yet been formulated quantitatively for quantum sensors. Furthermore, a general issue for quantum sensors arise in the multi-parameter Cramér-Rao inequality as the bounds are not tight and thus saturating them may not be achievable. Developing tighter bounds and strategies towards achieving them in many-body sensors require closer connections between quantum metrology and control theory. One important problem in multi-parameter sensing is the situation where Fisher information matrix is singular and thus sensing becomes impossible. Specifying the criteria which results in singularity of the Fisher information matrix in many-body systems and recipes to fix them requires further attention. Finally, experimental realization of quantum many-body sensors is a big objective which ultimately defines the success of these probes. Many physical platforms can implement many-body sensors, including nuclear magnetic resonance systems, Rydberg atoms, optical lattices, superconducting simulators, nitrogen vacancies and levitated objects. progress on these setups as well as new physical platforms open further opportunities for developing many-body quantum sensors with wide range of applications.
@@ Acknowledgements.
AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). VM acknowledges funding by the National Natural Science Foundation of China (Grants No. W2432005, No. 12374482, and No. 12050410251). RY acknowledges support from the National Natural Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242).

### v3

@@ XV.2 Machine Learning Based Approaches
The ongoing revolution in machine learning has already had a significant impact in the study of many-body quantum physics Carleo and Troyer (2017); Gray et al. (2018); Vicentini (2021). These range from identification of various exotic phases of matter in hitherto analytically challenging models Carrasquilla and Melko (2017); Greplova et al. (2020), to Hamiltonian learning Wang et al. (2017b), and investigation of foundational issues like Bell-nonlocality Deng (2018). Machine-learning algorithms, which learn from previously acquired data, also assist in solving optimization problems that arise in quantum sensing. For a general overview of machine-learning assisted quantum metrology, we refer the interested reader to the recent reviews by Huo et al Huo et al. (2022) and Huang et al Huang et al. (2024b). As we describe below, quantum sensing has already been shown as an especially fruitful area of application for machine learning methods.
@@ XV.2.2 Quantum machine learning strategies
XVI Experimental Realizations
The main aim of this article is to review the theoretical advancements of quantum many-body probes, as experimental developments have been explored in previous review articles, e.g. see Ref. Degen et al. (2017). Therefore, we keep this section very brief, mainly hinting the physical platforms in which quantum many-body probes are developed.
Quantum sensors have been developed across a diverse range of physical platforms, each harnessing unique properties of these systems. This diversity enables versatility in probe size, spanning atomic to microscales, and operational temperatures, ranging from ultra-cold atoms and ions to room temperature. These sensors have been used for detecting electric, magnetic, and gravitational fields, as well as measuring force, acceleration, and mass. Each platform offers distinct advantages, making them suitable for a wide array of sensing applications—from fundamental scientific research to practical technologies in navigation, medical imaging, and environmental monitoring. The key physical platforms for implementing quantum sensors include: ultra-cold atoms Kasevich and Chu (1992); Peters et al. (1999); Bongs et al. (2019); El-Neaj et al. (2020); Stray et al. (2022); Peters et al. (2001); Fixler et al. (2007); Tino (2021); Bloom et al. (2014); Hinkley et al. (2013); Takamoto et al. (2005), ion traps Leibfried et al. (2004); Maiwald et al. (2009); Biercuk et al. (2010); Sawyer et al. (2012); Brownnutt et al. (2015); Baumgart et al. (2016, 2016); Huntemann et al. (2016), atomic vapors Budker and Romalis (2007); Kominis et al. (2003); Dang et al. (2010); Balabas et al. (2010); Shah et al. (2007); Fernholz et al. (2008); Wasilewski et al. (2010); Xia et al. (2006), nuclear magnetic resonance devices Kitching et al. (2011); Fang and Qin (2012); Jiang et al. (2021a); Wang et al. (2022a); Wu et al. (2022); Jiang et al. (2024), solid state defects such as nitrogen vacancies in diamond Taylor et al. (2008); Clevenson et al. (2015); Jensen et al. (2014); Le Sage et al. (2012); Chernobrod and Berman (2005); Balasubramanian et al. (2008); Dolde et al. (2011); Zhao et al. (2012); Barson et al. (2017); Wu et al. (2021a); Vetter et al. (2022); Zhou et al. (2014); Xie et al. (2020); Schirhagl et al. (2014b); Patel et al. (2020), superconducting circuits Bal et al. (2012); Danilin et al. (2018); Wang et al. (2019a); Yu et al. (2025), photonic systems Holland and Burnett (1993); Mitchell et al. (2004); Pezzé et al. (2007); Higgins et al. (2007); Nagata et al. (2007); Ono et al. (2013); Xiao et al. (2024), and optomechanical devices Hu et al. (2024); Gavartin et al. (2012); Liu et al. (2022); Westerveld et al. (2021); Sansa et al. (2020); Li and Barker (2018); Fardian-Melamed et al. (2025); Fogliano et al. (2021); Pikovski et al. (2012); Chowdhury et al. (2023); Gosling et al. (2024); Forstner et al. (2012); McCormick et al. (2019); Aspelmeyer et al. (2014); Moser et al. (2013); Guzmán Cervantes et al. (2014); Krause et al. (2012); Chaste et al. (2012); Forstner et al. (2014); Degen et al. (2009); Rugar et al. (2004); Li et al. (2021); Liu et al. (2021a); Xia et al. (2023). For a comprehensive review on experimental implementation of quantum sensors, one can see Ref. Degen et al. (2017).
Thanks to recent advances on quantum technologies implementing quantum many-body probes are now viable. Quantum criticalities have been exploited for achieving quantum-enhanced sensitivity in several experiments. Nuclear Magnetic Resonance (NMR) systems have been exploited to harness quantum criticality with Heisenberg scaling Liu et al. (2021b) for sensing an external magnetic field. Quantum criticality in Rydberg atoms have also been exploited for sensing electric fields Ding et al. (2022). Furthermore, dipolar-coupled 13C nuclear spins in diamond have been employed to measure the frequency of time-varying periodic magnetic fields within a discrete time crystal framework Moon et al. (2024). By leveraging the time crystalline phase in these sensors, the probe’s lifetime can be exponentially enhanced by up to three orders of magnitude, enabling highly precise and selective frequency measurements. In a recent experiment, non-equilibrium quantum probes have also been realized in superconducting quantum processors to mitigate criticality slowing down mechanism and achieve Heisenberg scaling with respect to the encoding time Yu et al. (2025). All these experiments demonstrate the potential of quantum many-body systems to serve as quantum probes.
XVII Summary and Outlook
@@
Exploiting quantum features, such as entanglement, can enhance the precision of a sensor well-beyond the capacity of a classical sensor, the so-called quantum-enhanced sensing. Such enhanced in precision has placed quantum sensing as one of the pillars of emerging quantum technologies. In addition to such enhancement, the size versatility of quantum probes allows for achieving extremely high spatial resolution Ockeloen et al. (2013); Köse and Braun (2023). The potential applications are immense covering from biological monitoring Taylor and Bowen (2016) to mining Liu et al. (2024), atomic clocks Katori (2011) , navigation Baili et al. (2019) and space exploration Kohlrus et al. (2017).
@@
Interferometric quantum sensing has been the first demonstration of entanglement as a resource for sensing purposes. Some implementation of these types of sensors might be susceptible to decoherence as well as unwanted interaction between particles. An alternative approach is to exploit quantum many-body systems for sensing. Unlike interferometry-based quantum sensors, in many-body probes interaction plays a crucial role. In this review, we have explored different aspects of quantum sensing using many-body systems.
@@
Quantum many-body sensors have been used in two scenarios, namely equilibrium and non-equilibrium. In the former, quantum criticality has been identified as a resource for achieving quantum-enhanced sensitivity. There are several types of quantum criticalities, each with their own characteristics. However, in order to achieve quantum-enhanced sensitivity the criticality has to be accompanied by some sort of gap closing, for instance, Hamiltonian or Liouvillian spectral gap closing. This remarkable feature can be a useful guideline for identifying potential quantum sensors in new materials, which might be even designed by artificial intelligence Krenn et al. (2023). Although criticality-based quantum sensors offer several advantages, they also come with some drawbacks. First, the region over which quantum-enhanced precision is achievable can be very narrow, requiring probe’s fine tuning Montenegro et al. (2021). Second, preparing the probe near its critical point can be very resource-consuming because the preparation process slows down significantly as it approaches criticality Rams et al. (2018). In order to avoid these drawbacks, non-equilibrium quantum probes have also been proposed. In such systems quantum superposition and entanglement can be generated during the dynamics and thus complex initialization is not needed. Moreover, in non-equilibrium quantum probes, the evolution time acts as an additional parameter that can be used as a resource to enhance precision. In fact, the precision is often enhanced monotonically as time goes on Mishra and Bayat (2022). Nonetheless, unlike criticality-based sensors, in non-equilibrium probes the criteria for achieving quantum-enhanced sensitivity with respect to resources, e.g. system size, is not well-characterized. We summarize the advantages and disadvantages of each approach in Table 1.
@@
Despite current progress on many-body quantum sensing, there are still several open problems to explore. While criticality has been identified as a resource for sensing in equilibrium scenarios, the criteria which results in quantum-enhanced sensing is not specified in non-equilibrium probes. Another problem which requires further investigation is the performance of quantum sensors under imperfect situations, such as the presence of decoherence Matsuzaki et al. (2011). In general terms, while quantum sensors are naturally expected to be sensitive to the desired parameters, we would like them to be robust against other system’s parameters or noisy imperfections. Despite several works on noisy quantum metrology Escher et al. (2011b); Kołodyński and Demkowicz-Dobrzański (2013); Rossi et al. (2020); Demkowicz-Dobrzański and Maccone (2014); Demkowicz-Dobrzański et al. (2017); Dür et al. (2014); Nichols et al. (2016); Chaves et al. (2013); Alushi et al. (2024); Faist et al. (2023); Reiter et al. (2017); Mihailescu et al. (2023, 2024a), the notion of robustness has not yet been formulated quantitatively for quantum sensors. A related approach is the use of error-correction codes for quantum sensing Dür et al. (2014); Kessler et al. (2014); Zhou et al. (2018b); Unden et al. (2016); Zhou and Jiang (2020); Layden et al. (2019). For 1D interacting many-body systems, presence of even minimal disorder may induce Anderson localization Lahini et al. (2008), therefore such approaches may become extremely relevant for many-body sensing.
Furthermore, a general issue for quantum sensors arise in the multi-parameter Cramér-Rao inequality as the bounds are not tight and thus saturating them may not be achievable (Jiang et al., 2021c; Rubio and Dunningham, 2020; Fresco et al., 2022; Demkowicz-Dobrzański et al., 2020; Chen et al., 2022; Di Fresco et al., 2022; Liu et al., 2020; Albarelli et al., 2019; Mihailescu et al., 2024a; Yang et al., 2024b, 2025). Developing tighter bounds and strategies towards achieving them in many-body sensors require closer connections between quantum metrology and control theory. One important problem in multi-parameter sensing is the situation where Fisher information matrix is singular and thus sensing becomes impossible (Mihailescu et al., 2024a, 2025a; Yang et al., 2025; Mihailescu et al., 2025b; He and Paris, 2025; Frigerio and Paris, 2024; Bressanini et al., 2024; Mukhopadhyay et al., ming). Specifying the criteria which results in singularity of the Fisher information matrix in many-body systems and recipes to fix them requires further attention.
Another direction that holds immense promise is that of using multiple spatially separated sensors in a distributed manner, similar to classical sensor networks Rubio et al. (2020); Bringewatt et al. (2021); Proctor et al. (2018). So far, this work has mostly been performed for GHZ type Zang et al. (2024) and interferometric quantum sensors, see Ref. (Zhang and Zhuang, 2021) and references therein. However, very recently, proposals of distributed quantum sensing with many-body systems have also been put forward Hassani et al. (2025) focusing on privacy, i.e., individual sensor nodes having access to limited information as well as correlated noise estimation Brady et al. (2024).
Finally, experimental realization of quantum many-body sensors is a big objective which ultimately defines the success of these probes. Many physical platforms can implement many-body sensors, including nuclear magnetic resonance systems, Rydberg atoms, optical lattices, superconducting simulators, nitrogen vacancies and levitated objects. progress on these setups as well as new physical platforms open further opportunities for developing many-body quantum sensors with wide range of applications.
@@ Acknowledgements.
AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). VM acknowledges funding by the National Natural Science Foundation of China (Grants No. W2432005 and No. 12374482). RY acknowledges support from the National Natural Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242). SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012).